OpenStudy (anonymous):

Some stars are so far away that their position appear fixed as earth orbits the sun. Other stars, however, appear over time to shift their positions relative to the background of "fixed" stars. Suppose that the star shown below appears to shift through an arc of theta = 0 degrees 0 minutes and 1.5 seconds when viewed on the first day of winter and the first day of summer. If the distance from earth to the sun is about 1.5 x 10^8km, find the approximate distance from earth to the star.

OpenStudy (anonymous):

absolutely no idea :/

OpenStudy (anonymous):

I know how to do the actual problem, but my answer keeps ending up being .1 to large. I don't know where I'm making a mistake.

OpenStudy (superdavesuper):

If u show ur eqn, i will be happy to take a look :)

OpenStudy (superdavesuper):

But a difference of 0.1 is not too big right?

OpenStudy (anonymous):

The correct answer is 4.1 x 10^13 I keep ending up with 4.2 x 10^13

OpenStudy (anonymous):

This is the equation. 1.5 x 10^8 = 0.0004/360 (2)(pi)(r)

OpenStudy (superdavesuper):

Hm...going through it now.....1.5 seconds is 0.0004167....checking the rest now...

OpenStudy (superdavesuper):

I got 4.13x10^13 so it is 4.1x10^13 :)

OpenStudy (superdavesuper):

Think u shorten the 1.5 second value to 0.0004 causes the difference!

OpenStudy (anonymous):

Oh, okay. so I shouldn't round theta?

OpenStudy (superdavesuper):

U rounded it too early n too much - try it again n see if u get my results ;)

OpenStudy (anonymous):

I have no idea what I did wrong but I somehow ended up with 2.1 x 10^13....

OpenStudy (superdavesuper):

Oh sorry forgot to tell u because it is only from winter to summer, it is pi(r) - u got an extra 2x in ur eqn there

OpenStudy (anonymous):

Oh so I need to multiply 2.1 x 10^13 by 2?

OpenStudy (superdavesuper):

Yes but then...it will still be 4.2 instead of 4.1 :(

OpenStudy (superdavesuper):

What theat value u use in ur calc? i used 0.0004167 n got 4.1 as ans

OpenStudy (anonymous):

I used the same theta you did. But because 1.5 x 10^8 only applies to the first location of the earth, should I multiply that by 2 first and then solve?

OpenStudy (superdavesuper):

Well try that and see if u get 4.1 as answer :)

OpenStudy (anonymous):

I think I might be entering the information wrong. Every time I try the problem, I end up with different answers.

OpenStudy (superdavesuper):

Hahahaa want to see my steps?

OpenStudy (anonymous):

I got it! :D But can I please see them just to double check?

OpenStudy (superdavesuper):

Sure here it goes: theta=1.5/60/60=0.0004167 1.5x10^8 = theat/360 * pi * r = 3.636x10^-6 * r r = 4.13x10^13

OpenStudy (anonymous):

But the equation is S=theta/360 *2*pi*r What happened to the 2 in yours? Or am I just using the wrong equation?

OpenStudy (superdavesuper):

No, u are right...twice - ur eqn is right n u also right about doing 2x for 1.5x10^8 Same result though :)

OpenStudy (anonymous):

Okay, thank you for your help :)

OpenStudy (superdavesuper):

welcome n thanx 4 the medal!